# Finite-volume scattering on the left-hand cut

### v2

@@ 1 Introduction
Two other publications, of relevance for the left-hand cut in the finite volume context, have come to our attention during the completion of this work. These are ref. Sato:2007ms , concerning exponentially suppressed corrections to finite-volume scattering formula for $NN\to NN$, as well as ref. Meng:2021uhz , concerning a plane-wave basis for the quantization condition. The detailed relation of the formalism presented here to the earlier work is also a topic for future investigation.
@@ 2.3 Incorporating spin
$\big{[}iB\circ_{i\epsilon}iB-iB\circ_{{\sf pv}}iB\big{]}^{tv,t^{\prime}v^{\prime}}=\\ \frac{1}{2}\int\frac{d^{3}\bm{k}}{(2\pi)^{3}}iB^{tv,rs}(P,p,k)\frac{\pi\delta(E-\omega_{N}(\bm{k})-\omega_{N}(\bm{P}-\bm{k}))\,\delta_{rr^{\prime}}\delta_{ss^{\prime}}}{2\omega_{N}(\bm{k})2\omega_{N}(\bm{P}-\bm{k})}iB^{r^{\prime}s^{\prime},t^{\prime}v^{\prime}}(P,k,p^{\prime})\,,$ (54)
@@
${B}^{tv,rs}(P,p,k)$ $=\overline{u}_{t}^{\sigma}(\bm{p})\overline{u}_{v}^{\delta}(\bm{P}-\bm{p})B^{\sigma\delta,\alpha\beta}(P,p,k)u_{r}^{\alpha}(\bm{k})u_{s}^{\beta}(\bm{P}-\bm{k})$ (55)
@@
provided that on-shell momentum $k=(\omega_{N}(\bm{k}),\bm{k})$ is used on the left-hand side.
The key point here is that the spin indices appear through the Kronecker deltas in eq. (54). This translates to a corresponding modification to the phase-space factor, which can be written as
$\rho_{ss^{\prime}}(s)=\rho_{s_{1}s_{2},s_{1}^{\prime}s_{2}^{\prime}}(s)=\frac{k^{\star}_{\sf os}}{16\pi\sqrt{s}}\big{[}\delta_{s_{1}s_{1}^{\prime}}\delta_{s_{2}s_{2}^{\prime}}+\delta_{s_{1}s_{2}^{\prime}}\delta_{s_{2}s_{1}^{\prime}}\big{]}\,,$ (57)
where we use the indices $s$ and $s^{\prime}$ as a shorthand for the pairs of components, $s_{1}s_{2}$ and $s_{1}^{\prime}s_{2}^{\prime}$, respectively. Completing the derivation as above yields
$[\mathcal{M}]_{\ell s,\ell^{\prime}s^{\prime}}$ $=\frac{1}{[{\mathcal{K}^{-1}}]_{\ell s,\ell^{\prime}s^{\prime}}-i\rho_{ss^{\prime}}\,\delta_{\ell\ell^{\prime}}}\,,$ (58)
where
$\mathcal{K}_{\ell s,\ell^{\prime}s^{\prime}}(s)=16\pi\sqrt{s}\,\frac{\tan\delta_{\ell s,\ell^{\prime}s^{\prime}}(k)}{k}\bigg{|}_{k=k^{\star}_{\sf os}}\,,$ (59)
and $\delta_{\ell s,\ell^{\prime}s^{\prime}}(k)$ is the phase shift for two-to-two scattering with the spin components and orbital angular momentum as labelled. Crucially, a consequence of spin is that orbital angular momentum is no longer conserved so that two $\ell$ labels are needed.
A subsequent change of basis to total incoming and outgoing spin gives $\ell SM_{S},\ell^{\prime}S^{\prime}M_{S}^{\prime}$, and one can further combine these indices to label the amplitude in terms of total angular momentum. The full index set then takes the form: $JM_{J},\ell S,\ell^{\prime}S^{\prime}$. This is discussed in detail, for example in refs. Briceno:2014oea ; Briceno:2015csa and we do not give any more details here.
@@ 3.3 Reduction of the finite-volume two-particle loop
To continue the analysis of power-like $L$ dependence in $C_{L}(P)$ for $(2M_{N})^{2}<s<(2M_{N}+M_{\pi})^{2}$, we now analyze a single $NN$ loop, arising at any location of any skeleton expansion term, as shown in figure 5. We introduce generic functions to the left and right of the loop, denoted $\mathcal{L}(P,k)$ and *$\mathcal{R}^{*}(P,k)$ respectively. [^3] Here the $*{}^{*}$ on *$\mathcal{R}^{*}(P,k)$ indicates complex conjugation. This convention is useful when the functions represent the endcaps of a correlator constructed from an operator and its hermitian conjugate, as in eq. (60). In this case, $\mathcal{L}(P,k)=\mathcal{R}(P,k)$. These correspond to either endcap functions or Bethe-Salpeter kernels, and we omit other possible momentum arguments not relevant for the loop considered. The loop contribution can then be written as
@@
Additional contributions from the $k^{0}$ integral can be picked up from poles or branch cuts within $\mathcal{L}(P,k)$ and *$\mathcal{R}^{*}(P,k)$, but these lead to analytic functions of $\bm{k}$ that can be absorbed into the remainder term via $\mathcal{I}^{[1]}(P)\to\mathcal{I}^{[2]}(P)$. This holds for all objects that can appear in place of $\mathcal{L}(P,k)$ and *$\mathcal{R}^{*}(P,k)$, up to a caveat discussed in appendix A. The issue has to do with the distinction between $t$- and $u$-channel loops within kernels, defined according to the choice of momentum routing in the adjacent two-particle loops. The upshot is that it is always possible to find a routing for which $\mathcal{I}^{[2]}(P)$, $\mathcal{L}(P,k)$, and *$\mathcal{R}^{*}(P,k)$ have the desired analyticity and exponentially suppressed volume dependence. This also holds for the $t$- and $u$-channel single-pion exchanges (the problem with these first appears at a later step) and for objects containing lower left-hand cuts. The complete justification is given in section 3.6 below, together with appendix A.
Returning to eq. (74), we next decompose the functions $\mathcal{L}(P,k)$ and *$\mathcal{R}^{*}(P,k)$, with on-shell $k=(\omega_{N}(\bm{k}),\bm{k})$, into spherical harmonics with respect to the direction of the three-momentum in the CM frame:
@@
*$\mathcal{R}^{*}(P,k)\big{|}_{k^{0}=\omega_{N}(\bm{k})}$ $\equiv\sqrt{4\pi}\,Y^{*}_{\ell m}(\hat{\bm{k}}^{\star})|\bm{k}^{\star}|^{\ell}\,\widetilde{\mathcal{R}}^{*}_{\ell m}(P,|\bm{k}^{\star}|)\ ,$ (76)
@@ 3.6 Analytic structure of the Bethe-Salpeter kernel
$ig^{2}\mathcal{T}(P,p,p^{\prime})=-\frac{ig^{2}}{\big{(}\omega_{N}(\bm{p})-\omega_{N}(\bm{p}^{\prime})\big{)}^{2}-(\bm{p}-\bm{p}^{\prime})^{2}-M_{\pi}^{2}+i\epsilon}=-\frac{ig^{2}}{t-M_{\pi}^{2}+i\epsilon}\,,$ (99)
@@
The first key observation for eq. (99) is that it simply has no $s$ dependence, and therefore cannot generate any singularities in the complex $s$ plane. In addition to this, the momenta entering the $t$-channel exchange here guarantee that $t<0$ and thus that the singularity at $t=M_{\pi}^{2}$ is not encountered. This, in turn, implies that no power-like finite-volume effects arise in $C_{L}(P)$ due to this contribution. We will argue that this statement generalizes to any diagram that can be expressed as some number of $t$-channel exchanges.
@@ 3.7 Full decomposition of the finite-volume correlator
Here the ${}^{[1]}$ superscripts are used, as in section 3.3, to indicate intermediate quantities.
@@ 4.4 Analytic continuation
The momenta $p$, $p^{\prime}$, $P-p$ and $P-p^{\prime}$ can be put on shell and taken below threshold here. The key point is that the amplitude on the right-hand side is nested between two loops and thus its arguments are always real. Knowledge of $\mathcal{M}$ for real momenta can thus be exploited to obtain its subthreshold on-shell value.
@@ 5.3 $S$-wave dominance
In this final subsection, we consider the form of the quantization condition and the integral equations in the case that only the $S$-wave component of $\overline{\mathcal{K}}^{\sf os}$ is non-zero. Beginning with the quantization condition, eq. (169) becomes
@@ 6 Conclusions
As with much of the extensive literature relating finite-volume data to amplitudes, our derivations are based on a generic all-orders effective field theory of nucleons and pions with interactions only constrained by symmetries. Expressions are given for any spin of the nucleon, with emphasis on the toy spin-zero and the physical spin-half cases. We have made an effort to detail the steps that fail in the presence of subthreshold cuts, and to illustrate and motivate the adaptations required to solve the issue. To summarize, two key ingredients underpin our approach: the separation of the Bethe-Salpeter kernel into dangerous pion exchanges and a safe remainder, and the use of partial off-shellness as a tool for correctly representing the finite-volume correlator, in particular the fact that it does not contain left-hand cuts.

### v3

@@ 1 Introduction
Two other publications, of relevance for the left-hand cut in the finite-volume context, have come to our attention during the completion of this work. These are ref. Sato:2007ms , concerning exponentially suppressed corrections to finite-volume scattering formula for $NN\to NN$, as well as ref. Meng:2021uhz , which makes use of a plane-wave basis for the quantization condition. We compare our approach to that used in these earlier works in section 5.4.
@@ 2.3 Incorporating spin
$\big{[}iB\circ_{i\epsilon}iB-iB\circ_{{\sf pv}}iB\big{]}_{tv,t^{\prime}v^{\prime}}=\\ \frac{1}{2}\int\frac{d^{3}\bm{k}}{(2\pi)^{3}}iB_{tv,rs}(P,p,k)\frac{\pi\delta(E-\omega_{N}(\bm{k})-\omega_{N}(\bm{P}-\bm{k}))\,\delta_{rr^{\prime}}\delta_{ss^{\prime}}}{2\omega_{N}(\bm{k})2\omega_{N}(\bm{P}-\bm{k})}iB_{r^{\prime}s^{\prime},t^{\prime}v^{\prime}}(P,k,p^{\prime})\,,$ (54)
@@
${B}_{tv,rs}(P,p,k)$ $=\overline{u}_{t}^{\sigma}(\bm{p})\overline{u}_{v}^{\delta}(\bm{P}-\bm{p})B^{\sigma\delta,\alpha\beta}(P,p,k)u_{r}^{\alpha}(\bm{k})u_{s}^{\beta}(\bm{P}-\bm{k})\,,$ (55)
@@
provided that on-shell momentum $k=(\omega_{N}(\bm{k}),\bm{k})$ is used on the left-hand side. The sum runs over the two possible spin states, labeled here by the sign of the spin magnetic quantum number $r=\pm\frac{1}{2}$.
The key point here is that the spin-state indices appear through the Kronecker deltas in eq. (54). This translates to a corresponding modification to the phase-space factor, which can be written as
$\rho_{rs,r^{\prime}s^{\prime}}(s)=\frac{k^{\star}_{\sf os}}{16\pi\sqrt{s}}\,\delta_{rr^{\prime}}\delta_{ss^{\prime}}\,.$ (57)
Completing the derivation as above yields
~$\widetilde{\mathcal{M}}(s)$ $=\frac{1}{\widetilde{\mathcal{K}}(s)^{-1}-i\widetilde{\rho}(s)}\,,$ (58)
exactly matching eq. (24), but where all objects are now matrices in the combined orbital angular momentum and spin index space. The elements of these matrices are
$\widetilde{\rho}_{\ell mrs,\ell^{\prime}m^{\prime}r^{\prime}s^{\prime}}$ $=\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}(k^{\star}_{\sf os})^{2\ell}\rho_{rs,r^{\prime}s^{\prime}}\,,$ (59)
$\widetilde{\mathcal{K}}_{\ell mrs,\ell^{\prime}m^{\prime}r^{\prime}s^{\prime}}$ $=(k^{\star}_{\sf os})^{\ell+\ell^{\prime}}\,\mathcal{K}_{\ell mrs,\ell^{\prime}m^{\prime}r^{\prime}s^{\prime}}\,,$ (60)
~$\widetilde{\mathcal{M}}_{\ell mrs,\ell^{\prime}m^{\prime}r^{\prime}s^{\prime}}$ $=(k^{\star}_{\sf os})^{\ell+\ell^{\prime}}\,\mathcal{M}_{\ell mrs,\ell^{\prime}m^{\prime}r^{\prime}s^{\prime}}\,.$ (61)
Unlike the spinless case, we now need two sets of indices in the K-matrix and scattering amplitude because orbital angular momentum and spin orientation are not individually conserved and can thus change between incoming and outgoing states.
The elements $\mathcal{M}_{\ell mrs,\ell^{\prime}m^{\prime}r^{\prime}s^{\prime}}$ are projections of the amplitude to definite incoming and outgoing orbital angular momentum and single-particle spin states. We can write this projection as
$\mathcal{M}_{\ell mrs,\ell^{\prime}m^{\prime}r^{\prime}s^{\prime}}=\big{(}\langle\ell^{\prime}m^{\prime}|\otimes\langle s_{1}r^{\prime},s_{2}s^{\prime}|\,\big{)}\,\mathcal{M}\,\big{(}\,|\ell m\rangle\otimes|s_{1}r,s_{2}s\rangle\big{)}\,,$ (62)
where $|\ell m\rangle$ is a state of definite orbital angular momentum and $|s_{1}r,s_{2}s\rangle$ (with $s_{1}=s_{2}=\frac{1}{2}$ for nucleons) encodes the two-particle spin states. We can arrive at more useful set of indices by subsequent changes of basis, for example to definite total spin $\ell mSm_{S},\ell^{\prime}m^{\prime}S^{\prime}m_{S}^{\prime}$ via
$|\ell m,Sm_{S}\rangle$ $\equiv|\ell m\rangle\otimes|Sm_{S},s_{1}s_{2}\rangle$ (63)
$=\sum_{r,s}\big{(}\,|\ell m\rangle\otimes|s_{1}r,s_{2}s\rangle\big{)}\langle s_{1}r,s_{2}s|Sm_{S},s_{1}s_{2}\rangle\,,$ (64)
where $\langle s_{1}r,s_{2}s|s_{1}s_{2},Sm_{S}\rangle$ are the Clebsch-Gordon coefficients for the spin addition. One can further combine these indices to label the amplitude in terms of total angular momentum, using
$|Jm_{J},\ell S\rangle=\sum_{m,m_{S}}|\ell m,Sm_{S}\rangle\langle\ell m,Sm_{S}|Jm_{J},\ell S\rangle\,.$ (65)
The advantage is that total angular momentum is conserved, and the amplitude is diagonal in the indices $J,m_{J}$. The standard expressions for the K-matrix and amplitude in terms of the phase shift can be used with little alteration:
[…]
and
$\mathcal{M}_{Jm_{J}\ell S,J^{\prime}m_{J}^{\prime}\ell^{\prime}S^{\prime}}=\delta_{JJ^{\prime}}\delta_{m_{J}m_{J}^{\prime}}\mathcal{M}^{(J)}_{\ell S,\ell^{\prime}S^{\prime}}=\delta_{JJ^{\prime}}\delta_{m_{J}m_{J}^{\prime}}\frac{16\pi\sqrt{s}}{k^{\star}_{\sf os}\cot\delta^{(J)}_{\ell S,\ell^{\prime}S^{\prime}}(k^{\star}_{\sf os})-ik^{\star}_{\sf os}}\,,$ (67)
where the phase shifts $\delta^{(J)}_{\ell S,\ell^{\prime}S^{\prime}}$ are labelled by the total angular momentum and total incoming and outgoing orbital angular momentum and spin. For a more detailed discussion on the inclusion of spin, see for example refs. Briceno:2014oea ; Briceno:2015csa .
@@ 3.3 Reduction of the finite-volume two-particle loop
To continue the analysis of power-like $L$ dependence in $C_{L}(P)$ for $(2M_{N})^{2}<s<(2M_{N}+M_{\pi})^{2}$, we now analyze a single $NN$ loop, arising at any location of any skeleton expansion term, as shown in figure 5. We introduce generic functions to the left and right of the loop, denoted $\mathcal{L}(P,k)$ and $\mathcal{R}^{*}(P,k)$ respectively. [^3] Here the on $\mathcal{R}^{*}(P,k)$ indicates complex conjugation. This convention is useful when the functions represent the endcaps of a correlator constructed from an operator and its hermitian conjugate, as in eq. (68). In this case, $\mathcal{L}(P,k)=\mathcal{R}(P,k)$. These correspond to either endcap functions or Bethe-Salpeter kernels, and we omit other possible momentum arguments not relevant for the loop considered. The loop contribution can then be written as
@@
Additional contributions from the $k^{0}$ integral can be picked up from poles or branch cuts within $\mathcal{L}(P,k)$ and $\mathcal{R}^{*}(P,k)$, but these lead to analytic functions of $\bm{k}$ that can be absorbed into the remainder term via $\mathcal{I}^{[1]}(P)\to\mathcal{I}^{[2]}(P)$. This holds for all objects that can appear in place of $\mathcal{L}(P,k)$ and $\mathcal{R}^{*}(P,k)$, up to a caveat discussed in appendix A. The issue has to do with the distinction between $t$- and $u$-channel loops within kernels, defined according to the choice of momentum routing in the adjacent two-particle loops. The upshot is that it is always possible to find a routing for which $\mathcal{I}^{[2]}(P)$, $\mathcal{L}(P,k)$, and $\mathcal{R}^{*}(P,k)$ have the desired analyticity and exponentially suppressed volume dependence. This also holds for the $t$- and $u$-channel single-pion exchanges (the problem with these first appears at a later step) and for objects containing lower left-hand cuts. The complete justification is given in section 3.6 below, together with appendix A.
Returning to eq. (82), we next decompose the functions $\mathcal{L}(P,k)$ and $\mathcal{R}^{*}(P,k)$, with on-shell $k=(\omega_{N}(\bm{k}),\bm{k})$, into spherical harmonics with respect to the direction of the three-momentum in the CM frame:
@@
$\mathcal{R}^{*}(P,k)\big{|}_{k^{0}=\omega_{N}(\bm{k})}$ $\equiv\sqrt{4\pi}\,Y^{*}_{\ell m}(\hat{\bm{k}}^{\star})|\bm{k}^{\star}|^{\ell}\,\widetilde{\mathcal{R}}^{*}_{\ell m}(P,|\bm{k}^{\star}|)\ ,$ (84)
@@ 3.6 Analytic structure of the Bethe-Salpeter kernel
$ig^{2}\mathcal{T}(P,p,p^{\prime})=-\frac{ig^{2}}{\big{(}\omega_{N}(\bm{p})-\omega_{N}(\bm{p}^{\prime})\big{)}^{2}-(\bm{p}-\bm{p}^{\prime})^{2}-M_{\pi}^{2}+i\epsilon}\,,$ (107)
@@
The first key observation for eq. (107) is that it has no $s$ dependence, by virtue of the fact that it is evaluated at $p=(\omega_{N}(\bm{p}),\bm{p})$ and $p^{\prime}=(\omega_{N}(\bm{p}^{\prime}),\bm{p}^{\prime})$, with the magnitude of $\bm{p}$ and $\bm{p}^{\prime}$ left independent of $E$ (by allowing $(E-\omega_{N}(p),\bm{P}-\bm{p})$ and its primed counterpart to be off shell). In this sense it cannot generate any singularities in the complex $s$ plane. In addition to this, the momenta entering the $t$-channel exchange here guarantee that $t<0$ and thus that the singularity at $t=M_{\pi}^{2}$ is not encountered. This, in turn, implies that no power-like finite-volume effects arise in $C_{L}(P)$ due to this contribution. We will argue that this statement generalizes to any diagram that can be expressed as some number of $t$-channel exchanges.
@@ 3.7 Full decomposition of the finite-volume correlator
Here the [1] superscripts are used, as in section 3.3, to indicate intermediate quantities.
@@
We emphasize again that the object we are projecting here is not the on-shell exchange containing the sub-threshold pole leading to the left-hand cut. As remarked at the end of section 2.2, such an expansion is expected to have poor convergence on the cut. Instead, what we project is the partially off-shell exchange of eq. (118) (cf. the Case 2 kinematics described in the previous section). This object does not contain a pole for real $|\bm{p}^{\star}|$, $|\bm{p}^{\prime\star}|$, and can therefore be expanded in partial waves without convergence problems.
@@ 4.4 Analytic continuation
The momenta $p$, $p^{\prime}$, $P-p$ and $P-p^{\prime}$ can be put on shell and taken below threshold here. The key point is that the amplitude on the right-hand side is nested between two loops and thus its arguments are always real. Knowledge of $\mathcal{M}$ for real momenta can thus be exploited to obtain its subthreshold on-shell value. These methods can similarly be applied to analytically continue the amplitude into the complex plane. Whether using the method above or directly trying to solve the integral equations for complex momenta, it is important to account for the singularities contained in the amplitude and $\mathcal{K}^{\mathcal{T}}$. Depending on the specific external momenta considered, these can cross the path of integration on the real line, requiring contour deformation.
@@ 5.3 $S$-wave dominance
In this subsection, we consider the form of the quantization condition and the integral equations in the case that only the $S$-wave component of $\overline{\mathcal{K}}^{\sf os}$ is non-zero. Beginning with the quantization condition, eq. (175) becomes
@@
5.4 Comparison to previous work
In this subsection, we briefly compare our results with those of refs. Sato:2007ms ; Meng:2021uhz .
We begin with ref. Sato:2007ms , one of the first papers to emphasize the importance of single-pion exchange in exponentially suppressed volume corrections to the Lüscher quantization condition. The aim of the publication is to point out that such effects may be large enough to bias the extracted phase shifts for sufficiently small volumes, and to use non-relativistic nuclear chiral effective field theory (EFT) to numerically estimate the size of the corrections. In contrast to our formalism, ref. Sato:2007ms uses a version of the Bethe-Salpeter kernel that is given solely by the single-pion exchange. This is referred to as a potential and denoted by $V$. Comparing eq. (32) of the earlier work with our eq. (127), one can see that our two approaches are brought closer by setting $\overline{B}=0$ in our derivation and taking the non-relativistic expansion.
The additional volume effects identified in ref. Sato:2007ms then correspond to the sum-integral-difference acting on a difference between the on-shell and off-shell versions of $V$, contained in a quantity called $\mathbb{F}$, defined in eq. (20) of ref. Sato:2007ms . The authors assume that the finite-volume energies are not on the left-hand cut, so $\mathbb{F}$ is real-valued, and this leads to real-valued exponentially suppressed corrections to the standard Lüscher scattering formalism. Other differences from our approach include the neglect of partial-wave mixing in the finite volume, the restriction to interactions described by the specific EFT, and the fact that the former work restricts attention to (and provides numerical estimates for) the two-nucleon system with physical masses.
The more recent work of ref. Meng:2021uhz advocates using a plane-wave basis, as opposed to the partial-wave basis typically used for two-particle quantization conditions. Also here, the authors suggest using an EFT-based framework to describe the one-pion exchange, and show that applying this in the plane-wave basis circumvents the partial-wave mixing effects by providing an alternative approach to treat the broken rotational symmetry of the finite volume.
The main difference between our approach and that of ref. Meng:2021uhz is that our quantization condition [given in eq. (130)] is defined over a combined plane-wave (indexed by $\bm{k}^{\star},\bm{k}^{\prime\star}$) and angular momentum (indexed by $\ell m,\ell^{\prime}m^{\prime}$) basis. Our alternative, equivalent form of the quantization condition [given in eq. (175)] is defined by a determinant only over the angular momentum space. Importantly, both versions of our main result depend on the same K-matrix analog, the quantity we call $\overline{\mathcal{K}}^{\sf os(\ell)}(s)$, and this does not depend on the magnitudes or directions of the plane-wave basis indices $\bm{k}^{\star}$ and $\bm{k}^{\prime\star}$. Implementing our method in practice will require the use of various models or parameterizations of $\overline{\mathcal{K}}^{\sf os(\ell)}(s)$, which will include tests to ensure that a sufficient set of partial wave components are included so that the neglected higher partial waves are numerically irrelevant. Such tests have already been performed in ref. Meng:2021uhz , and it will be instructive to quantitatively compare the results.
The above summary concerns qualitative similarities and differences between the methods, and a direct numerical test and quantitative comparison is still outstanding and urgently needed. This is left to future work, as is a more detailed theoretical exploration of regimes in which the formulas are equivalent.
@@ 6 Conclusions
As with much of the extensive literature relating finite-volume data to amplitudes, our derivations are based on a generic all-orders EFT of nucleons and pions with interactions only constrained by symmetries. Expressions are given for any spin of the nucleon, with emphasis on the toy spin-zero and the physical spin-half cases. We have made an effort to detail the steps that fail in the presence of subthreshold cuts, and to illustrate and motivate the adaptations required to solve the issue. To summarize, two key ingredients underpin our approach: the separation of the Bethe-Salpeter kernel into dangerous pion exchanges and a safe remainder, and the use of partial off-shellness as a tool for correctly representing the finite-volume correlator, in particular the fact that it does not contain left-hand cuts.